7.Iterative methods of the solutions of systems of equations are ________________
a.finite
b.sequential
c.infinite
d.non-sequential

8.Stirling\'s formula for interpolation is given by
a.\\(P_n (x) = f (x_0) + \\frac{s}{2}[df_\\frac{1}{2} + d_\\frac{-1}{2}] + \\frac{s^2}{2}d^2f_0\\)
b.\\(P_n (x) = f (x_0) + \\frac{s}{2}[df_\\frac{1}{2} + d_\\frac{-1}{2}] + \\frac{s^2}{2!}d^2f_0\\)
c.\\(P_n (x) = f (x_0) + \\frac{s - 1}{2}[df_1+ d_\\frac{-1}{2}] + \\frac{s^2}{2!}d^2f_0\\)
d.\\(P_n (x) = f (x_0) + \\frac{s}{2}[df_\\frac{1}{2} + d_\\frac{-1}{2}] + \\frac{s^2}{2!}df_0\\)
1
Expert's answer
2019-09-12T08:57:07-0400

7.Iterative methods of the solutions of systems of equations are infinite


Answer: c.infinite


8.Stirling\'s formula for interpolation is given by 


"P_n (x) = f (x_0) + \frac{s}{2}[df_\frac{1}{2} + d_\frac{-1}{2}] + \frac{s^2}{2!}d^2f_0"

Answer: b.P_n (x) = f (x_0) + \frac{s}{2}[df_\frac{1}{2} + d_\frac{-1}{2}] + \frac{s^2}{2!}d^2f_0